OpenStudy (anonymous):

Divide (10x4y3 + 5x3y2 - 15x2y - 25x2y4) by -5x2y

hero (hero):

\[\frac{10x^4y^3 + 5x^3y^2 - 15x^2y - 25x^2y^4}{-5x^2y} = -\frac{10x^4y^3 + 5x^3y^2 - 15x^2y - 25x^2y^4}{5x^2y} \] \(=-\left(\dfrac{10x^4y^3 + 5x^3y^2 - 15x^2y - 25x^2y^4}{5x^2y}\right) \) \(=-\left(\dfrac{10x^4y^3}{5x^2y} + \dfrac{5x^3y^2}{5x^2y} - \dfrac{15x^2y}{5x^2y} - \dfrac{25x^2y^4}{5x^2y}\right)\)

hero (hero):

That should help you out a bit